a) Find the linear and quadratic approximations of

sin(0.96π) tan(0.26π) + (0.96)^2 (0.26)

b) Compare these approximations with the actual value.


1
Expert's answer
2021-09-23T00:35:22-0400

(a) Given function is: "f(x)=sin(0.96\u03c0) tan(0.26\u03c0) + (0.96)^2 (0.26)"

Since no point is given so we are calculating the value at point 0.

The linear approximation to "f(x)=\sin \left(\frac{\pi}{25}\right) \tan \left(\frac{13 \pi}{50}\right)+\frac{3744}{15625} \ at \ x_{0}=0 \\"

A linear approximation is given by: "L(x) \approx f\left(x_{0}\right)+f^{\prime}\left(x_{0}\right)\left(x-x_{0}\right) ."

We are given that "x_{0}=0 ."

Firstly, find the value of the function at the given point:

"y_{0}=f\left(x_{0}\right)=\sin \left(\frac{\pi}{25}\right) \tan \left(\frac{13 \pi}{50}\right)+\frac{3744}{15625}"

Secondly, find the derivative of the function, evaluated at the point: "f^{\prime}(0) ."

Next, evaluate the derivative at the given point to find slope.

"f^{\prime}(0)=0"

Plugging the values found, we get that: "L(x) \approx \sin \left(\frac{\pi}{25}\right) \tan \left(\frac{13 \pi}{50}\right)+\frac{3744}{15625}+0(x-(0)) ."

Or, more simply: "L(x) \approx \sin \left(\frac{\pi}{25}\right) \tan \left(\frac{13 \pi}{50}\right)+\frac{3744}{15625} ."

Answer: "L(x) \approx \sin \left(\frac{\pi}{25}\right) \tan \left(\frac{13 \pi}{50}\right)+\frac{3744}{15625} \approx 0.373082337744149"


A quadratic approximation is given by: "Q(x) \approx f\left(x_{0}\right)+f^{\prime}\left(x_{0}\right)\left(x-x_{0}\right)+\frac{1}{2} f^{\prime \prime}\left(x_{0}\right)\left(x-x_{0}\right)^{2} ."

We are given that "x_{0}=0 ."

Firstly, find the value of the function at the given point:

"y_{0}=f\left(x_{0}\right)=\sin \left(\frac{\pi}{25}\right) \tan \left(\frac{13 \pi}{50}\right)+\frac{3744}{15625}"

Secondly, find the derivative of the function, evaluated at the point: "f^{\prime}(0) ."

Find the derivative: "f^{\prime}(x)=0"

Next, evaluate the derivative at the given point.

"f^{\prime}(0)=0"

Now, find the second derivative of the function evaluated at the point: "f^{\prime \prime}(0) ."

Find the second derivative: "f^{\prime \prime}(x)=0"

Next, evaluate the second derivative at the given point.

"f^{\prime \prime}(0)=0 ."

Plugging the found values, we get that:

"Q(x) \approx \sin \left(\frac{\pi}{25}\right) \tan \left(\frac{13 \pi}{50}\right)+\frac{3744}{15625}+0(x-(0))+\frac{1}{2}(0)(x-(0))^{2}\\
 
Simplify: Q(x) \approx \sin \left(\frac{\pi}{25}\right) \tan \left(\frac{13 \pi}{50}\right)+\frac{3744}{15625} .\\
Answer: Q(x) \approx \sin \left(\frac{\pi}{25}\right) \tan \left(\frac{13 \pi}{50}\right)+\frac{3744}{15625}"


(b) Actual Value of function is: "\sin \left(\frac{\pi}{25}\right) \tan \left(\frac{13 \pi}{50}\right)+\frac{3744}{15625}"

So, we can say that actual value of function is same as linear and quadratic approximation's value because no variable is there.